# LibriQuote: A Speech Dataset of Fictional Character Utterances for Expressive Zero-Shot Speech Synthesis

### v1

@@ Abstract
Text-to-speech (TTS) systems have recently achieved more expressive and natural speech synthesis by scaling to large speech datasets. However, the proportion of expressive speech in such large-scale corpora is often unclear. Besides, existing expressive speech corpora are typically smaller in scale and primarily used for benchmarking TTS systems. In this paper, we introduce the LibriQuote dataset, an English corpus derived from read audiobooks, designed for both fine-tuning and benchmarking expressive zero-shot TTS system. The training dataset includes 12.7K hours of read, non-expressive speech and 5.3K hours of mostly expressive speech drawn from character quotations. Each utterance in the expressive subset is supplemented with the context in which it was written, along with pseudo-labels of speech verbs and adverbs used to describe the quotation (e.g. “he whispered softly”). Additionally, we provide a challenging 7.5 hour test set intended for benchmarking TTS systems: given a neutral reference speech as input, we evaluate system’s ability to synthesize an expressive utterance while preserving reference timbre. We validate qualitatively the test set by showing that it covers a wide range of emotions compared to non-expressive speech, along with various accents. Extensive subjective and objective evaluations show that fine-tuning a baseline TTS system on LibriQuote significantly improves its synthesized speech intelligibility, and that recent systems fail to synthesize speech as expressive and natural as the ground-truth utterances. The dataset and evaluation code are freely available. Audio samples can be found at https://libriquote.github.io/.
Code — https://github.com/deezer/libriquote
Dataset — https://huggingface.co/datasets/gasmichel/LibriQuote
Introduction
@@
Recently, text-to-speech (TTS) systems have significantly improved by scaling to large-scale, multi-domain speech corpora (Shen et al. 2023; Wang et al. 2023; Kharitonov et al. 2023; Jiang et al. 2023a; Anastassiou et al. 2024; Wang et al. 2024; Chen et al. 2024; Wang et al. 2025), exhibiting remarkable speech naturalness, quality and voice following abilities. These large-scale corpora typically contain speech derived from audiobooks (Kahn et al. 2019; Pratap et al. 2020; Kang et al. 2024) or in-the-wild data (He et al. 2024). While in-the-wild speech may include diverse forms of expressivity and speaking style, it is sometimes believed that audiobook speech corpora may lack such expressiveness (He et al. 2024). In contrast, we claim that human reading in audiobooks, especially readings of fictional works, offer extensive prosodic variability, as they often involve an alternation between neutral, non-expressive speech and expressive, sometimes emotional, utterances. Neutral speech usually corresponds in a book to narration parts, these story fragments enunciated by the narrator – may it be a first or third person view, or free indirect speech – while expressive speech is employed during dialogues between fictional characters. Dialogue is an essential narrative function (Genette 1983) that allows fictional character to portray themselves, expressing aspects of their personality, emotional states, and their own representation of the fictional world. As the story unfolds, characters navigate dynamically through their own emotion arc (Vishnubhotla et al. 2024), which is often portrayed during audiobook reading. Indeed, numerous blog posts dedicated to amateurs and written by professional narrators emphasize the importance of impersonation, or the act of bringing life to fictional characters (Gonzalez 2021; Browm 2021; Goodwin 2024).
In this work, we put fictional characters at the forefront, and propose LibriQuote, a large-scale dataset of expressive utterances derived from LibriVox111https://librivox.org recordings. Notably, LibriQuote differs from standard large-scale audiobook datasets (Pratap et al. 2020; Kahn et al. 2019; Kang et al. 2024) by segmenting speech utterances in two ways: character utterances that involve expressive and diverse spech, and narration parts that correspond to a neutral, formal reading style. In addition, we provide as side information the context in which each character utterance appears, supplemented with pseudo-labels of potential speech verbs and adverbs used to describe the utterance (e.g. “he whispered softly”), automatically annotated and further validated by humans. These pseudo-labels provide rich information that we further exploit to construct an expressive data subset. Through an extensive alignment pipeline involving audio preparation and transcription from LibriVox, text matching from Project Gutenberg222https://www.gutenberg.org/ and quotation detection (Bamman, Underwood, and Smith 2014), we collected more than 3 million quotations uttered by fictional characters that we systematically aligned with their book counterpart, resulting in nearly 5400 hours of speech. An overview of the pipeline is provided in Figure 1. Qualitative analysis of the data reveals that LibriQuote contains a large set of emotions and accents, facilitating the training and benchmarking of expressive TTS systems, as depicted in Figure 2.
Our objective and subjective experiments demonstrate the effectiveness of LibriQuote. When evaluated on our test set, modern open-source TTS systems fail to synthesize utterances as expressive and as natural as the ground-truth quotations, suggesting that further research should be conducted to build more expressive systems. When used as a fine-tuning corpora, LibriQuote-train significantly improves speech intelligibility of a baseline TTS system. Besides, we derive an expressive data subset using pseudo-labels of speech verbs and adverbs, which we show is a data-efficient alternative to the full LibriQuote-train set.
Our contribution are as follows:
We release publicly LibriQuote, a dataset derived from audiobooks dedicated to zero-shot TTS. LibriQuote comes with standard train/dev/test splits. LibriQuote-train comprises 5.3K hours of training quotations, as well as 12.7K hours of training narration segments, while LibriQuote-test contains 7.5 hours of quotations distributed among 15 speakers.
We release publicly the code and evaluation samples used in objective and subjective experiments to encourage benchmarking on LibriQuote-test.
We show through extensive objective and subjective evaluations that modern TTS systems fail to synthesize utterances as expressive and as natural as in real audiobooks.
We show that fine-tuning a baseline TTS system on LibriQuote-train significantly improves speech intelligibility on external audiobooks and in-the-wild benchmarks.
Existing Datasets
We briefly review the composition and characteristics of existing datasets used for zero-shot speech synthesis.
Audiobook-based Speech Datasets
Audiobooks have traditionally been the de-facto open-source data available for training TTS systems. The LibriVox project records thousands of audiobooks derived from public-domain books, read by volunteered speakers. LibriSpeech (Panayotov et al. 2015), a 1000-hours corpus of English speech data extracted from LibriVox that covers 2400 speakers, has been extensively used in the literature to train and benchmark Automatic Speech Recognition (ASR) and TTS systems. LibriTTS (Zen et al. 2019) provides an enhanced version of LibriSpeech, with better audio quality and utterance segmentation, but supports a smaller amount of training data, and LibriSpeech-PC (Meister et al. 2023) restored the punctuation and capitalization of LibriSpeech audio transcripts. Libri-Light (Kahn et al. 2019) is the largest English audiobook corpus of open data, with approximately 60,000 hours of English speech derived from LibriVox, but does not provide audio transcriptions, making it unpractical to train supervised ASR or classic TTS systems. The MLS datasest (Pratap et al. 2020) is composed of 50,000 hours of multilingual audiobooks and proposes audio transcriptions that were aligned with the original text. LibriHeavy (Kang et al. 2024) improved Libri-Light by aligning audio segments derived from recordings and their text counterpart in the book from which the segment was taken, but do not provide book contextual information in the vicinity of audio segments. These 30-seconds audio segments are created at sentence boundaries, thus they may contain multiple quotations and narration parts, alternating between neutral (narration) and expressive (quotation) speech. Because of this alternation patterns, complex pitch distributions arise in each training instance, which may hinder the learning of TTS systems. In contrast, we extract speech from a subset of LibriVox, Fiction Novels, separating training instances between neutral (narration) and expressive (quotation) segments, preserving contextual information around quotations. In fiction, quotations from fictional characters is an essential communicative medium between the reader and the characters, revealing their emotions and personality through dialogue.
Expressive Speech Datasets
@@
Audiobook narration typically alternates between neutral, narrative speech and expressive speech during character acting. Székely et al. (2012) design a system to capture expressive speech utterances from audiobooks, and validate their approach on a single Hungarian audiobook. To the best of our knowledge, no work have focused on capturing expressive audio segments in audiobooks at a large scale. However, several efforts towards the creation of expressive read speech corpora have been made. Emotional speech datasets (Busso et al. 2008; Cao et al. 2014; Livingstone and Russo 2018; Zhou et al. 2021) that assign discrete emotion labels to speech utterances have been extensively used to evaluate emotional synthesize of TTS systems (Jiang et al. 2023b; Wang et al. 2024, 2025). The L2-ARCTIC corpus (Zhao et al. 2018) proposes an 11 hour corpus with ten speakers, spanning 5 different accents. EXPRESSO (Nguyen et al. 2023) introduces an high-quality dataset of read and improvised speech across 26 expressive styles. In LibriQuote, we create audio segments that match character utterances and which are thus likely to be more expressive than narration parts. Besides, we extract narrative information indicating how an utterance should be spoken, providing a natural language pseudo-label of expressivity.
LibriQuote Dataset
This section describes the steps involved in preparing LibriQuote, followed by descriptive statistics. Then we present our methodology to extract narrative information for each utterance. Figure 1 provides an overview of the full process.
Dataset Construction
@@
We mainly follow the recipe of LibriHeavy (Kang et al. 2024) to align audio files downloaded from LibriVox to their original text. However, we filter out books that are unlikely to contain fictional characters, such as bibliographical work or other type of non-fictional works. We describe the filtering and alignment stages below.
Audio Preparation
@@
We start by browsing the LibriVox API333https://librivox.org/api/info to collect recording information. We only extract English recordings corresponding to Fiction books. This step is necessary to filter out non-fiction books that are unlikely to contain fictional characters. Each recording is downsampled to 16kHz for further processing, but we provide links to download the original audio files as well, allowing LibriQuote to support systems operating on higher audio quality. Note that LibriVox also contains Dramatic Readings, which are recordings where each fictional character is acted by a unique speaker. We also applied the alignment process described below to these recordings, but do not include them in LibriQuote as they would require an additional speaker diarization step. We plan to release these alignments in future works.
Text Preparation
@@
For each audio recording, we download the corresponding book text from Project Gutenberg. Some books have different text providers, but we found by manual inspection that automatically downloading from these websites was either too complicated, or resulted in texts with very poor quality often due to OCR errors. In such cases, we browse Project Gutenberg to search for the target book, and discard it if not available on the platform. We then perform quotation detection using BookNLP444https://github.com/booknlp/booknlp/blob/main/booknlp/english/litbank˙quote.py on each book. We discard every book containing less than quotations, as we found that these book were typically showing quotation detection errors. For other books, we keep quotation positions in the original text for further alignments.
@@
Audio Transcription
We start by segmenting each audio recording into 30-second segments, with a 2 second overlap at each side. These segments are then transcribed with a Zipformer-Transducer ASR model trained on LibriSpeech555https://github.com/k2-fsa/icefall/pull/1058, and combined by leveraging word-level timestamps to obtain a full timestamped transcript of the recording.
Text-Audio Alignment
Since audio recordings often involve a full chapter, this step is necessary to find the corresponding chapter text within the full original book, and to have fine-grained alignment at the word level. This alignment process involves two stages. In the first stage, close matches between a transcription and the book text are constructed. These close matches correspond to pairs where is an index in the transcription and an index in the book text. Then, a coarse alignment is produced by taking the longest chain of pairs such that and . The second stage produces a final alignment by concatenating Levenshtein alignments (Lcvenshtcin 1966) between the transcription and the text segment produced by the longest chain of pairs. We refer the reader to Kang et al. (2024) for additional information regarding the alignment process.
Quotation Alignment
The output of the text-audio alignment is an array of words in the original text, timestamped in the audio recording. We leverage the timestamps and the output of the quotation detection system to segment each audio into a set of read quotations. Additionally, we build audio segments from narration paragraphs (i.e. parts that are not quotations) in each book. The final construction stage involved assigning contextual information around quotations. Based on the alignments, we match each quotation with a window of around words occurring before and after it, with boundaries defined by paragraphs. With LibriQuote, we release the preprocessed texts as well as full paragraph and quotation alignments. Thus, it is possible to derive a contextual window of any desirable length around quotations.
Descriptive Statistics
We display the main characteristics of LibriQuote in Table 1.
@@
The filtering and alignment process resulted in a total of 12723 training hours of narration and 5359 training hours of quotation across 2991 unique books and 3314 unique speakers. As expected, narration parts constitute the majority of audiobook recording in terms of duration, but quotations still represent almost 50% of the total number of training utterances. Indeed, the average audio segment duration is 11.8 seconds for narration parts and 5.5 seconds for quotations. Figure 3 displays the histogram of utterance durations. LibriVox does not provide gender information of its readers, so the balance between male and female speaker in the training set is unclear.
| Verbs | Adverbs | |||||
| Support | 344 | 56 | ||||
| P | R | F | P | R | F | |
| Phi-4 | 0.85 | 0.93 | 0.89 | 0.74 | 0.81 | 0.77 |
| Phi-4 | 0.92 | 0.86 | 0.89 | 0.95 | 0.61 | 0.74 |
@@
We designed the development and test sets to reflect the following challenge in expressive TTS: given a reference timbre from neutral speech and the text to generate, can we synthesize a quotation with the correct level of expressiveness? Each quotation serves as the expressive speech, and we match them with the nearest narration utterance from the book context that acts as the neutral reference speech. Without the context, the task can be particularly hard as the quotation text itself may be misleading with regard to the true emotion, which may rather be specified in the context. As an example, the exclamation mark in the quotation “It’s so good!” he whispered softly might suggest it shall be uttered with a loud voice, but the narrative information clarifies that it is actually being whispered. Therefore, we also provide contextual information in the vicinity of quotations to encourage further research on predicting expressivity from textual context for audiobook synthesis.
The development dataset contains randomly selected utterances from speakers and books in the train set. It thus does not encompass zero-shot speech synthesis, but aims at evaluating expressive synthesis from neutral speech for seen speakers. The test set contains 8 male and 7 female unseen speakers, for a total of 7.5 quotation hours. We ensure no speaker overlap with LibriSpeech train and dev speakers.
@@
To further analyze LibriQuote test set, we explore its emotional and accent variability. Following Wang et al. (2024), we compute emotion representations using Emotion2Vec666https://huggingface.co/emotion2vec/emotion2vec˙plus˙base (Ma et al. 2024) and accent representations using an ECAPA-TDNN model trained on Common Voice777https://huggingface.co/Jzuluaga/accent-id-commonaccent˙ecapa (Ardila et al. 2020; Zuluaga-Gomez et al. 2023). We display t-SNE projections of emotion representations for LibriQuote-test quotations and their reference narration utterance in Figure 2. For comparison, we also display t-SNE projections of 5734 utterances randomly drawn from all LibriHeavy segments that contain at least one quotation. We see that LibriQuote-test quotations convey a larger set of predicted emotions compared to their narration reference and LibriHeavy segments: only 67% of quotations are predicted as neutral against 87% and 91% for the reference and LibriHeavy respectively. Besides, we found that accent representations are well clustered, indicating a broad coverage of English accents. Further details can be found in the Technical Appendix.
Narrative Information Extraction
| WER | SIM-O | FPC | MCD | A-Sim | E-Sim | CMOS | MOS | |
| Ground Truth | 7.2 | - | - | - | 0.0 | 3.58 | ||
| SparkTTS | 5.5 | 0.45 | 0.65 | 9.1 | 0.61 | 0.69 | -0.97 | 3.26 |
| F5-TTS | 8.1 | 0.52 | 0.63 | 9.5 | 0.66 | 0.70 | -0.98 | 3.47 |
| MaskGCT | 9.3 | 0.55 | 0.63 | 9.5 | 0.65 | 0.72 | -0.98 | 3.36 |
Distinguishing between ”story” (what is happening) and ”discourse” (how it is told) is a key aspect of narrative understanding. NLP research has extensively focused on various aspects of narrative discourse, as theorized by Genette (1983). Recently, a variety of works explored the usage of Large Language Models (LLMs) to model elements of narrative understanding, such as dialogue (Michel et al. 2025), character roles and information extraction (Stammbach, Antoniak, and Ash 2022; Gurung and Lapata 2024), narrativity (Hatzel and Biemann 2024) or structure (Soni et al. 2023).
The current work is grounded in the perspectival issues of fictional characters, including point of view and dialogues, anchored in the voice linking function of Genette’s theoretical framework. In fiction, utterances are sometimes supplemented with narrative information in the form of speech verbs and adverbs (e.g. “he whispered softly”). These indicators have been shown to produce positive effects during reading, such as enhancing reading comprehension (Wolters, Kim, and Szura 2022) or character identification (Van Krieken, Hoeken, and Sanders 2017). For audiobook narrators, they are a valuable resource to set the tone and point of view of the character. Other contextual information such as the emotional context of a character are likely to also impact speech delivery, but are often prone to subjective interpretation and are harder to automatically infer (Vishnubhotla et al. 2024). Thus, by extracting these narrative cues and filtering out less expressive utterances (e.g. “she said”), we can derive a high-quality subset that spans a broad prosodic range.
Method

### v2

# Computational Narrative Understanding for Expressive Text-to-Speech

@@ Abstract
Recent advances in text-to-speech (TTS) have been driven by large, multi-domain speech corpora, yet the expressive potential of audiobook data remains underexamined. We argue that human-narrated audiobooks, particularly fictional works, contain rich and diverse prosodic cues arising from the natural alternation between neutral narration and expressive character dialogue. Building from this observation, we introduce LibriQuote, a large-scale 5.3K hours of expressive speech drawn from character quotations. Each quote is supplemented with contextual pseudo-labels for speech verbs and adverbs that characterize the intended delivery of direct speech (e.g., “he whispered softly”). We found that fine-tuning a flow-matching model on LibriQuote yields substantial improvements in expressivity and intelligibility, while training from scratch enhances expressiveness of an autoregressive TTS model. Benchmarking on LibriQuote-test highlights significant variability across systems in generating expressive speech. We publicly release the dataset, code, and evaluation resources to facilitate reproducibility. Audio samples can be found at https://libriquote.github.io/.
Computational Narrative Understanding for Expressive Text-to-Speech
Gaspard Michel†∗ gmichel@deezer.com Elena V. Epure†⋄ elena.epure@idiap.ch
Christophe Cerisara∗ christophe.cerisara@loria.fr
† Deezer Research, Paris, France ∗ Loria, Nancy, France ⋄ IDIAP, Martigny, Switzerland
1 Introduction
Recently, text-to-speech (TTS) systems have significantly improved by scaling to larger, multi-domain speech corpora Shen et al. (2023); Chen et al. (2025); Kharitonov et al. (2023); Jiang et al. (2024); Anastassiou et al. (2024); Wang et al. (2024); Chen et al. (2024); Wang et al. (2025), exhibiting remarkable naturalness, quality and voice following abilities. Typically, these large-scale corpora contain speech derived from audiobooks Kahn et al. (2019); Pratap et al. (2020); Kang et al. (2024) or in-the-wild data He et al. (2024). While the latter may include diverse expressivity and speaking styles, it is sometimes claimed that audiobooks may lack such expressiveness He et al. (2024).
@@ 1 Introduction
Contrary to this, we argue that human readings of audiobooks, especially fictional works, offer extensive prosodic variability through a natural alternation between neutral speech and expressive utterances. Neutral speech usually corresponds to narrative parts, the fragments enunciated by the narrator, while expressive speech can be employed during dialogues between fictional characters. Dialogue is an essential narrative function Genette (1983) that allows fictional characters to portray themselves and express aspects of their personality, emotional states, and world representation. As the story unfolds, characters navigate dynamically through their own emotion arc Vishnubhotla et al. (2024), which can be portrayed during audiobook reading. Indeed, numerous blog posts written by professional narrators emphasize the importance of impersonation, or the act of bringing life to fictional characters Gonzalez (2021); Browm (2021); Goodwin (2024).
In this work, we put fictional characters at the forefront, by creating a large-scale dataset of expressive quotes and proving its effectiveness for expressive TTS. Our dataset, LibriQuote, differs from standard large-scale audiobook datasets Pratap et al. (2020); Kahn et al. (2019); Kang et al. (2024) by relying on a narrative-aware segmentation focused on character quotes that exhibit expressivity and diversity, as opposed to narration, which corresponds to a more neutral reading style. We automatically annotate, and validate with humans, speech cues specific to quotations as contextual information: pseudo-labels of speech verbs and adverbs used in the narrative to characterize direct speech utterances (e.g. “he whispered softly”). LibriQuote-train includes 3300 speakers and 5.3K hours of quotations and 12.7K hours of narration, while LibriQuote-test provides 7.4 hours of quotations from 15 speakers.
Our findings are the following:
1.
Fine-tuning a state-of-the-art autoregressive TTS system with LibriQuote-train does not directly improves expressivity, but yields substantial improvements in speech intelligibility on in-domain and out-of-domain data (Section 4). Besides, when training from scratch the same TTS system with LibriQuote-train, we increase speech expressivity, but at the cost of reduced intelligibility due to substantially reduced training data Ye et al. (2025). In contrast, fine-tuning a flow-matching model directly enhances both speech expressivity and intelligibility. Finally, we show that using both narrations and quotations contained in LibriQuote-train is a promising direction for training expressive and intelligible TTS systems.
2.
The benchmarking of various TTS systems on LibriQuote-test reveals large discrepancies in TTS systems’ ability to synthesize expressive speech (Section 5). This indicates that LibriQuote-test represents a challenging benchmark that could foster advances in developing expressive TTS.
The dataset is available at the following links:
Huggingface: https://huggingface.co/datasets/gasmichel/LibriQuote
2 Background and Motivation
2.1 Related Speech Datasets
Audiobooks have long been the de-facto open-source data for training TTS systems. The LibriVox project records thousands of public-domain books read by volunteer speakers. LibriSpeech Panayotov et al. (2015), a 1000-hour English corpus covering 2400 speakers and extracted from LibriVox, has been widely used to train and benchmark TTS systems. LibriTTS Zen et al. (2019) offers an enhanced version of LibriSpeech with better audio quality and utterance segmentation, though with less training data, while LibriSpeech-PC Meister et al. (2023) restores punctuation and capitalization in LibriSpeech transcripts.
Libri-Light Kahn et al. (2019) is the largest open English audiobook corpus at about 60,000 hours of LibriVox speech but lacks audio transcriptions. The MLS dataset Pratap et al. (2020) includes 50,000 hours of multilingual audiobooks and provides transcriptions aligned with the original text. LibriHeavy Kang et al. (2024) improves Libri-Light by aligning audio segments with their corresponding book text, but does not provide contextual book information around segments. These 30-second segments are cut at sentence boundaries and may include multiple quotations and narration, alternating between neutral (narration) and expressive (quotation) speech.
@@ 2.1 Related Speech Datasets
Moreover, several efforts towards the creation of expressive read speech corpora have been made. Emotional speech datasets Busso et al. (2008); Cao et al. (2014); Livingstone and Russo (2018); Zhou et al. (2021) that assign discrete emotion labels to speech utterances have been extensively used to evaluate emotional synthesize of TTS systems Jiang et al. (2024); Wang et al. (2024, 2025). The L2-ARCTIC corpus Zhao et al. (2018) proposes an 11 hour corpus with ten speakers, spanning 5 different accents. EXPRESSO Nguyen et al. (2023) introduces an high-quality dataset of read and improvised speech across 26 expressive styles.
2.2 Gaps in Current Datasets
Prior works that use audiobooks as a resource for speech data, such as LibriTTS or LibriHeavy, have overlooked aspects of narrative discourse. They either completely discard character quotations, or combine them with other quotations and/or neutral narrative parts in their audio segments, which are constructed from random sentence breaks. Such segments are thus likely to exhibit multiple "distinct" pitch distributions, possibly alternating between neutral, emotional and/or pitch-varying speech. We illustrate this effect in Figure 1.
Specifically, we performed quotation detection on 120000 segments from LibriHeavy-small, and calculated their pitch standard deviation in the corresponding audio segments. We found that 75% of segments are only narrative parts, while the remaining 25% contain from 1 to 12 quotations. We see that pitch standard deviation typically increases with the number of quotations within the segment (Spearman , ).
Modeling these complex pitch distributions is a challenging task. It can hinder the training of TTS systems, which may end up focusing on learning to model the simpler, narrative parts instead. By shifting the focus from formal narration to character quotations, LibriQuote provides varied direct speech samples that span a broader range of emotions, at scale. These traits are desirable for generating high-quality, human-like speech He et al. (2024). Moreover, we supplement these quotations with narrative information indicating how an utterance should be spoken, providing a natural language pseudo-label for expressivity.
3 The LibriQuote Dataset
Distinguishing between "story" (what is happening) and "discourse" (how it is told) is a key aspect of narrative understanding. NLP research has extensively focused on various aspects of narrative discourse, as theorized by Genette (1983). Recently, a variety of works explored the usage of Large Language Models (LLMs) to model elements of narrative understanding, such as dialogue Michel et al. (2025), character roles and information extraction Stammbach et al. (2022); Gurung and Lapata (2024), narrativity Hatzel and Biemann (2024) or structure Soni et al. (2023).
The current work is grounded in the perspectival issues of fictional characters, including point of view and dialogues, anchored in the voice linking function of Genette’s theoretical framework. In fiction, utterances are sometimes supplemented with narrative information in the form of speech verbs and adverbs (e.g. “he whispered softly”). These indicators have been shown to produce positive effects during reading, such as enhancing reading comprehension Wolters et al. (2022) or character identification Van Krieken et al. (2017). For audiobook narrators, they are a valuable resource to set the tone and point of view of the character. Other contextual information such as the emotional context of a character are likely to also impact speech delivery, but are often prone to subjective interpretation and are harder to automatically infer Vishnubhotla et al. (2024).
3.1 Data Creation Pipeline
@@
Our data creation pipeline is similar to the one proposed for LibriHeavy Kang et al. (2024). However, we filter out books that are unlikely to contain fictional characters, such as bibliographical or other types of non-fictional works. We detail the filtering and audio-text alignment stages below.
3.1.1 Audio Preparation
@@
We start by browsing the LibriVox API111https://librivox.org/api/info to collect recording information. We only extract English recordings corresponding to Fiction books. This step is necessary to filter out non-fiction books that are unlikely to contain fictional characters. Each recording is downsampled to 16kHz for further processing, but we provide links to download the original audio files as well, allowing LibriQuote to support systems operating on higher audio quality. LibriVox also includes Dramatic Readings, where each fictional character is performed by a different speaker. We exclude these recordings from LibriQuote, as they would require an additional speaker diarization step.222We applied the same alignment process to these recordings and plan to release these alignments in future work.
3.1.2 Text Preparation
@@
For each audio recording, we download the corresponding book text from Project Gutenberg. Some books have external text providers, but manual inspection revealed that automatically downloading from these websites was often too complicated or yielded low-quality texts, frequently due to OCR errors. In such cases, we turned to Project Gutenberg to search for the target book and discarded it if it was not available on the platform. We then perform quotation detection using BookNLP333https://github.com/booknlp/booknlp/blob/main/booknlp/english/litbank_quote.py on each book. We discard every book containing less than quotations, as we found that these books typically yield quotation detection errors. For other books, we keep quotation positions in the original text for further alignments.
@@
3.1.3 Audio Transcription
We start by segmenting each audio recording into 30-second segments, with a 2 second overlap at each side. These segments are then transcribed with a Zipformer-Transducer ASR model trained on LibriSpeech444https://github.com/k2-fsa/icefall/pull/1058, and combined by leveraging word-level timestamps to obtain a full timestamped transcript of the recording.
3.1.4 Text-Audio Alignment
Since audio recordings often involve a full chapter, this step is necessary to align the corresponding chapter text within the full original book. This alignment process involves two stages. In the first stage, close matches between a transcription and the book text are constructed. These close matches correspond to pairs where is an index in the transcription and an index in the book text. Then, a coarse alignment is produced by taking the longest chain of pairs such that and . The second stage produces a final alignment by concatenating Levenshtein alignments Lcvenshtcin (1966) between the transcription and the text segment produced by the longest chain of pairs. We refer the reader to Kang et al. (2024) for additional information regarding the alignment process.
3.1.5 Quotation Alignment
The output of the text-audio alignment is an array of words in the original text, timestamped in the audio recording. We leverage the timestamps and the output of the quotation detection system to segment each audio into a set of read quotations. Additionally, we build audio segments from narration paragraphs (i.e. parts that are not quotations) in each book, and trimmed long start end silences for each audio segment. The final construction stage involved assigning contextual information around quotations. Based on the alignments, we match each quotation with a window of around words occurring before and after it, with boundaries defined by paragraphs. With LibriQuote, we release the preprocessed texts as well as full paragraph and quotation alignments. Thus, it is possible to derive a contextual window of any desirable length around quotations.
3.2 Dataset Characteristics
Table 1 shows LibriQuote’s descriptive statistics.
@@ 3.2 Dataset Characteristics
The filtering and alignment process resulted in a total of 12723 training hours of narration and 5359 training hours of quotation across 2991 unique books and 3314 unique speakers. As expected, narration parts constitute the majority of audiobook recording in terms of duration, but quotations still represent almost 50% of the total number of training utterances. Indeed, the average audio segment duration is 11.8 seconds for narration parts and 5.5 seconds for quotations. LibriVox does not provide gender information of its readers, so the balance between male and female speaker in the training set is unclear.
@@
We designed the dev and test sets to evaluate whether synthesized speech in quotations have the correct level of expressiveness. Each quotation serves as the expressive speech, and we match them with the nearest narration utterance from the book context that acts as the neutral reference speech. Synthesizing these quotations can be particularly challenging, as the text alone may be misleading with regard to the true emotion, which is often specified by the surrounding context. As an example, the exclamation mark in the quotation “It’s so good!” he whispered softly might suggest it shall be uttered with a loud voice, but the narrative information clarifies that it is actually being whispered. Thus, we also provide contextual information in the vicinity of quotations to encourage further research on predicting expressivity from textual cues for audiobook TTS. The test set contains 8 male and 7 female unseen speakers, for a total of 7.5 hours.
The dev set contains random utterances from speakers and books in the train set. It thus does not encompass zero-shot speech synthesis, but aims at evaluating expressive synthesis from neutral speech for seen speakers. However, we ensure no speaker-overlap with the test set.
@@
We start the exploration of LibriQuote-test by computing Word Error Rate (WER) whith Whisper-Large v3 and an automated metric for speech naturalness, UTMOS Saeki et al. (2022), ranging from 1 (unnatural) to 5 (very natural). We found that utterances exhibiting high WER often contain only a few words and proper names such as “Hey, Jean”. We thus decided to remove 205 utterances from the test set that contain less than two words. Other segments with high WER were sometimes spoken with a strong accent, which might affect Whisper transcription, but we decided to keep these segments. UTMOS analysis reveal an average value of 3.48, where most lower valued utterances occur when a quotation is spoken with very high emotion intensity. We thus chose not to filer out these utterances, as they correspond to desired expressivity.
@@
To further analyse LibriQuote test set, we explore which emotions are conveyed in quotations and reference narrations. Following Wang et al. (2024), we compute emotion representations using Emotion2Vec555https://huggingface.co/emotion2vec/emotion2vec_plus_base Ma et al. (2024). We found that LibriQuote-test quotations convey a larger set of predicted emotions compared to their narration reference and LibriHeavy segments: only 67% of quotations are predicted as neutral against 87% and 91% for the reference and LibriHeavy respectively. Further details can be found in Appendix A.
3.3 Contextual Cues for Direct Speech
We extract narrative cues indicating the speaking style of a quotation (described by the narrator), using a contextual window spanning one paragraph before and after the quotation. For this, we rely on several LLMs as they have shown promising results in literary dialogue understanding Piper and Bagga (2024); Michel et al. (2025).